day, 28 July 99

Marj Corcoran (corcoran@physics.rice.edu)
Wed, 28 Jul 1999 15:48:30 -0500 (CDT)

Shift Summary for DAY shift, 28-July-99

On shift: S. Glazov, M. Corcoran

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 8.3
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: tape change, a few kfrends
==================================================================

Runs ended this shift:
Run 14045 E832_beam, ended just as the shift ended

Tapes brought to FCC: QK*182-184, QKB062

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:

Smooth E832 data-taking.

Elliott occasionally warns of inefficient HCC bits, but it goes away by the
next spill. A few KTEVANA errors, fixed by kfrend.

At very end of shift, MI problems, est. 20 mins for beam

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14045 245 9:30 SDAQ errors, several processes were not running,
but apparently they restarted on their own.

14045 9:45 Bud comes into counting room to do some drilling;
installing some panels on the console

14045 333 11:30 Beam goes away.

14045 338 11:35 Beam returns.

14045 404-406 13:05 Trigger in error, kfrend. Then KTEVANA errors,
second kfrend cures problem

14045 417 13:20 KTEVANA errors again, kfrend again cures

14045 470 14:35 Notice SEM is low, around 2E12. We have had about
10 spills at lower intensity

14045 492 15:15 Tapes full, end run

14046 15:30 Tape drive tps3d1 on ktev1 loaded a new tape as
write protected, even though it wasn't. We reloaded
the tape, all OK, but had to drt_fresh_start

14046 15:30 MI problems just as we try to start new run.

14046 4 15:40 Beam returns, all seems OK. Evening crew arrives.

===================================================================

Followup notes on previous shift

===================================================================
Items to follow up on next shift:

He tubes will be changed tomorrow. Pressure will be a bit low, but it will
last through the night, so don't panic. Or, if you do panic, page Dave N.